### v1

@@ 3 Results
Under Scenario 6 (with RD1 chosen) and Scenario 7 (no SBSP selected, as RD2 is not adopted by the model), we compare weekly average power outputs for SBSP, solar, wind, and hydro (Figure 7). Throughout the year in Scenario 6, RD1 consistently delivers 300–350 GW with only minor dips around the spring and autumn equinoxes. By contrast, solar, wind, and hydro in both scenarios hover below 200 GW on average and exhibit notable seasonal swings—solar peaks in summer, while wind and hydro tend to peak in winter or other off-peak solar seasons.
@@
With RD1’s inclusion (Scenario 6), wind generation declines by approximately 50%, while solar output drops by 70–80%, illustrating RD1’s displacement of intermittent renewables. In contrast, Scenario 7—effectively a no-SBSP scenario—maintains a mix of variable renewables throughout the year, which require more storage to maintain the supply-demand balance.
@@
Figure 7: Weekly Power Output Variability of SBSP and Other Renewables under Scenario 6-7 in 2050. Panels (a) and (b) present a comparative analysis of RD1 and RD2(No-SBSP), respectively, with Solar, Wind, and Hydro in terms of power output and variability. Orange represents Solar, blue represents Wind, green represents Hydropower, and red represents SBSP (RD1 or RD2). Solid lines indicate weekly average power output, while shaded bands denote the range of weekly maximum and minimum power fluctuations.
Under Scenarios 6 and 7, we observe marked shifts in energy storage behavior (Fig. 8). In Scenario 7, RD2 is not selected by the model and the system effectively operates without SBSP, relying on pumped hydro (orange), batteries (blue), and hydrogen (green) to meet demand. Here, pumped hydro exhibits moderate fluctuations, while batteries show frequent charge–discharge cycles with pronounced seasonal swings, and hydrogen undergoes longer charging periods. The total storage usage (red line) displays wide fluctuations throughout the year, underscoring the need for intensive storage interventions in the absence of SBSP.
@@
By contrast, in Scenario 6, RD1 is deployed and profoundly alters storage utilization patterns. During summer, the average power drawn from storage drops by more than 50% relative to Scenario 7, thanks to RD1’s stable high-output supply. This is reflected in narrower deviation bands for all storage technologies, indicating a more consistent operating regime. In winter, however, total weekly storage requirements increase slightly, revealing a continued need for long-duration seasonal solutions. Although battery usage declines dramatically (by 78% annually), hydrogen storage takes on a larger role in winter, highlighting how RD1’s near-baseload profile mitigates short-term imbalances without fully eliminating seasonal challenges.
Figure 8: Weekly Power Variability and Charging/Discharging Dynamics of Storage Technologies under Scenario 6 and 7 in 2050. This figure illustrates the weekly variability in total power for Pumped Hydro Storage (PHS), Battery, and Hydrogen storage, showing both charging (positive) and discharging (negative) phases, across Europe under Scenario 6 and 7. Solid lines indicate weekly average power, while shaded bands denote the range of weekly maximum and minimum power fluctuations.
We evaluated the regional distribution of SBSP capacity for RD1 across three 2050 scenarios (Fig. 9), each reflecting different terrestrial-renewable cost assumptions—labeled “Maximum,” “Average,” and “Minimum.” This range spans pessimistic to optimistic outlooks for terrestrial wind and solar costs.
In all three scenarios, RD1 consistently demonstrates substantial total capacity and broad geographic coverage, with deployment in 24-27 countries, representing 73-82% of the modeled region. Germany, the United Kingdom, and France show the largest installed capacities (darkest shading). Although the number of countries deploying RD1 is not significantly affected by different cost assumptions, the total installed capacity undergoes a substantial change, decreasing from 444.6 GW in Scenario 3 to 372.8 GW in Scenario 6 and 281.3 GW in Scenario 9. Since SBSP can provide power to multiple regions simultaneously, as long as the orbital generation capacity meets the required output, it has the potential to supply electricity to the majority of countries within the modeled region.
Figure 9: Regional Distribution of RD1 Installed Capacity Across Cost Scenarios in 2050. This figure shows the installed capacity of RD1 under scenarios (denoted as S3, S6, and S9) in 2050.
Fig. 10 illustrates the changes in cross-border transmission capacity under nine scenarios (S2-S10), categorized by maximum renewable energy costs (left column: Scenarios 2/4, 3), average costs (middle column: Scenarios 5/7, 6), and minimum costs (right column: Scenarios 8/10, 9). In each column, the top panel shows the results with no SBSP (same with RD2), and the bottom panel with RD1.
Comparing no SBSP (top row) to RD1 (bottom row), we observe that RD1 consistently reduces the need for interregional lines across all cost scenarios. This reflects RD1’s higher annual availability, which can supplant remote generation and reduce reliance on cross-border imports or exports. Specifically, the total transmission capacity for HVAC decreases by 27-28% in all three scenarios, while the total capacity for HVDC decreases by 31-70%. This indicates that RD1 is particularly effective in reducing long-distance DC transmission. The large variation in HVDC capacity changes across different scenarios suggests that RD1 is more competitive when renewable energy costs are higher, leading to greater deployment and, consequently, larger reductions in transmission capacity requirements.
Overall, lower-cost scenarios (right column) exhibit more pronounced transmission expansions (with thicker and/or more lines), whereas higher-cost scenarios (left column) require less new cross-border capacity. In scenarios with minimum renewable energy costs, certain regions specialize heavily in low-cost generation, stimulating larger interregional flows to countries where generation is relatively more expensive, thus driving greater transmission investment. In contrast, when renewables are costly (left column), the system relies more on alternative resources—such as local SBSP or other dispatchable options—thereby reducing the economic incentive for extensive cross-border trade.
Figure 10: Regional Changes in Transmission Capacity in 2050. This figure illustrates the total transmission capacity changes across Scenario 2 to 10 (denoted as S2 to S10) for 2050. HVAC denotes high-voltage alternating current transmission lines, and HVDC denotes high-voltage direct current transmission lines.
Figure 11 illustrates the variation in storage capacities for pumped hydro (orange), battery (purple), and hydrogen (green) across Europe under nine different 2050 scenarios (S2–S10). In scenarios incorporating RD1 (bottom row), several Northern European countries, including Norway, Finland, and Sweden, experience a significant reduction in total storage capacity compared to the no-SBSP case (top row). In certain cases, this reduction is almost complete for specific technologies; for instance, in Norway and Sweden, the pumped hydro storage capacity can decrease by as much as 95%, while in Finland, hydrogen storage capacity is reduced to zero. These results suggest that RD1’s high annual availability significantly reduces the need for local storage capacity. Hydrogen storage capacity expands notably under RD1 in many parts of Europe, with significant increases observed in countries such as Germany, the Netherlands, Italy, and the Czech Republic.
Figure 11: Regional Changes in Storage Capacity in 2050. This figure illustrates the total storage capacity changes, including PHS, Battery, and Hydrogen, across Scenario 2 to 10 (denoted as S2 to S10) for 2050.
Figure 12: Changes in Installed Capacity in Response to SBSP Capital Costs in 2050. The panel (a) and (b) compares the installed capacities of terrestrial solar, wind, and SBSP (RD1 and RD2) under varying SBSP capital costs. The x-axis represents the capital cost of SBSP in EUR/kW, while the y-axis represents the corresponding installed capacity (GW).
Sensitivity analysis in 2050 scenarios - Scenarios 11-26
Building on our 2050 scenario findings that rely on NASA’s cost projections, we next conduct a sensitivity analysis to reflect the uncertainty of these estimates and to identify plausible cost thresholds for SBSP compared to terrestrial renewables (i.e., solar). Specifically, we adjust SBSP capital costs by up to ±65% of RD1’s baseline values—100–450 EUR/kW—representing both optimistic and more pessimistic trajectories. This range ensures the net-zero role exploration that at the upper end, SBSP remains non-competitive, while at the lower end, it can substantially displace other generation options. Throughout these runs, the costs of other technologies remain fixed at their 2050 estimates (“Average”).
@@
Fig. 12 shows that once SBSP capital costs fall to around 14× or 9× the 2050 solar PV cost in the RD1 and RD2 cases respectively, SBSP starts displacing wind and solar in a complementary role. A further  35% cost reduction beyond these thresholds enables SBSP to emerge as the dominant baseload technology, surpassing conventional renewables in capacity deployment. However, at these levels, SBSP may produce surplus generation (due to reduced load factors, see SI Section 4), effectively meeting both baseload demand and balancing short- and long-duration fluctuations. However, current prohibitive costs of SBSP are at least 1-2 orders of magnitude higher than these complementary or baseload cost benchmarks.
@@
The results further highlight the shifting role of long-duration storage as SBSP scales from small to large deployment (Fig. 13). Initially, as SBSP capital costs decline to approximately 14× (RD1) or 9× (RD2) the projected 2050 solar PV cost, the capacity of short- and medium-duration storage technologies, such as batteries and PHS, steadily decreases. Hydrogen storage follows a similar trend, with a sharp reduction occurring when SBSP costs reach around 11–14× (RD1) or 8–9× (RD2) the 2050 solar PV benchmark. In this phase, SBSP effectively serves as a seasonal supplement, partially displacing the need for hydrogen storage.
@@
However, once SBSP capital costs fall below approximately 11× (RD1) or 8× (RD2) the 2050 solar PV cost, hydrogen storage capacity surges dramatically, reaching up to 178% (RD1) and 174% (RD2) of its original level in the absence of SBSP. This sudden expansion indicates that at these cost levels, winter electricity demand imbalances drive extensive reliance on hydrogen storage, despite SBSP’s ability to manage short-term fluctuations. Indeed, while total storage capacity increases during this phase, the actual energy supply from storage continues to decline (Fig. 14). In this context, hydrogen provides up to 96% of its total annual supply during winter, highlighting the persistent need for long-duration storage to complement SBSP, even when SBSP reaches cost-competitive levels.
@@
A sharp decline in total storage capacity occurs when SBSP capital costs drop further to 8× (RD1) or 5× (RD2) the 2050 solar PV cost, marking the point at which SBSP surplus capacity becomes dominant. This shift signals SBSP’s ability to provide continuous baseload power, significantly reducing the system’s reliance on both short- and long-duration storage solutions.
@@
Figure 13: Impact of SBSP Capital Cost on Optimized Energy Storage Capacity in 2050. The panel (a) and (b) illustrates the relationship between SBSP’s capital cost and the optimized energy storage capacity of three storage technologies: battery, hydrogen, and pumped hydro storage (PHS), as well as their total capacity, in the 2050 scenario. The energy storage capacity represents the maximum amount of electricity that storage devices can deliver when fully charged.
@@
Figure 14 illustrates the transition of SBSP into the dominant electricity source as its capital cost decreases. When SBSP capital costs reach approximately 9× (RD1) or 6× (RD2) the projected 2050 cost of solar PV, SBSP surpasses other technologies in electricity generation. A further decline to around 3× the solar PV cost results in RD1 and RD2 contributing 99% and 98% of total electricity generation, respectively.
Notably, RD1 and RD2 exhibit opposing impacts on energy storage requirements. As RD1 deployment increases, total storage supply declines, shrinking to just 9% of its initial level. In contrast, RD2—due to its greater output variability—necessitates additional storage to compensate for fluctuations. Despite these divergent trends in overall storage supply, the share of battery consistently increases across both scenarios. When SBSP capital costs fall to 3× the cost of solar PV, electricity supply from battery accounts for 100% of total stored energy in both RD1 and RD2 cases. This finding underscores a critical threshold: if SBSP costs decline sufficiently, a power system without long-term storage could become viable.
Fig. 15 shows that the total annual system cost declines significantly as SBSP capital costs decrease, once RD1 or RD2 is deployed. When their capital costs fall to 3× the projected 2050 solar PV cost, total system costs are reduced by 52% for RD1 and 33% for RD2, underscoring SBSP’s substantial potential to lower overall energy expenditures. Throughout this transition, SBSP’s total generation cost initially rises due to increased deployment but subsequently declines sharply as generation costs decrease, ultimately displacing biomass and wind to become the dominant contributor to generation costs. Additionally, the reduction in SBSP generation costs leads to a corresponding decline in transmission costs, with DC transmission experiencing the most pronounced decrease. This finding further supports the conclusion from Fig. 10, which highlights SBSP’s superior ability to reduce long-distance DC transmission requirements.
@@
Figure 14: Impact of SBSP Capital Cost on Annual Electricity Supply in 2050. The panel (a) and (b) illustrate the changes in total annual electricity supply from all generation and storage technologies under varying RD1’s and RD2’s capital costs in 2050, respectively. The numbers above each bar indicate the total annual electricity supply from the three storage technologies combined, while the black dashed line represents the annual electricity demand.
Figure 15: Impact of SBSP Capital Cost on Total Annual Costs of Systems in 2050. The panel (a) and (b) illustrate the changes in total annual costs of whole systems under varying RD1’s and RD2’s capital costs in 2050, respectively. The numbers above each bar indicate the total annual costs of different scenarios.
@@ 4 Discussions
This study explores the system-wide impacts of two representative SBSP designs—RD1 (near-constant, lower TRL) and RD2 (partially intermittent, higher TRL)—under diverse European energy system conditions. When we apply NASA’s 2050 cost and performance assumptions, RD1 consistently emerges as a cost-competitive option, capable of reducing total system costs by 7–15% relative to a no-SBSP scenario. Its near-baseload profile not only displaces a substantial share of wind and solar but also cuts battery storage usage by as much as 78%. At the same time, long-duration solutions, especially hydrogen, still expand in winter months, confirming SBSP’s ability to resolve short-term imbalances while underscoring ongoing seasonal needs. In contrast, RD2 never appears in the least-cost mix at NASA’s 2050 cost projections, implying that its improved but still partially intermittent orbital geometry and higher capital expenses render it uncompetitive against both RD1 and terrestrial renewables. Notably, the stable, cross-network generation offered by RD1’s power supply markedly reduces long-distance DC transmission by 31–70% across regions, suggesting that a centrally coordinated, high-availability resource can diminish reliance on interregional electricity flows.
@@
Our sensitivity analysis pinpoints the capital-cost thresholds at which each design moves from non-competitive to supplementary, and ultimately to a dominant baseload resource. At present, SBSP’s costs are 1–2 orders of magnitude above these break-even points. As RD1 and RD2 costs approach 9–14× that of 2050 solar PV, SBSP begins to complement terrestrial renewables; dropping further to around 6–9× enables a baseload role with surplus capacity for balancing. Below these thresholds, SBSP could start to become the primary generation source, though the system might still reply on energy storage. Conversely, RD2—given its partially intermittent orbit—would also require cost reductions to at least match the lower complementary threshold, plus a reliance on short-duration storage to handle variability.
@@
Collectively, our findings show that near-constant SBSP (as exemplified by heliostat-based RD1) may deliver significant system benefits, but also depends on technological breakthroughs (e.g., advanced orbital assembly) to overcome its currently low TRL. In the shorter term, less-intermittent planar-based design RD2, with higher TRL, may make it an easier candidate for early demonstration, but only if its capital costs are reduced sufficiently to remain economically viable alongside terrestrial solar, wind, and energy storage. A coordinated development strategy could thus focus on RD2 demonstrations first to refine core SBSP technologies—such as wireless power transmission and modular in-orbit assembly—while concurrently accelerating R&D for SBSP designs with more continuous power generation. Such a plan balances the practical realization of a tangible full-scale SBSP prototype with the longer-term potential for near-baseload SBSP to disrupt the broader renewables landscape once its more complex capabilities achieve higher maturity.
@@
To achieve the cost-competitiveness of SBSP, achieving a gigawatt-scale SBSP platform by the 2030s requires several key breakthrough advances. Although space-qualified photovoltaics can achieve high efficiency, deploying vast arrays—potentially on the order of several or tens of km² in a single platform—remains unproven. In-orbit manufacturing and autonomous assembly, both at low to mid TRL, need to advance in parallel with large-scale orbital robotics and modular construction. Equally crucial is wireless power transmission. While small-scale demonstrations, such as Caltech’s recent low-power orbit-to-Earth test [10], validate core concepts, maintaining and scaled-up operating highly focused microwave beams over tens of thousands of kilometers needs to be proven. Consequently, our capacity-expansion modeling, though primarily focused on the system-level economic and dispatch implications of SBSP, must treat the 2030 horizon as an early full-scale demonstration phase, with true commercialization more plausibly materializing before 2050 if all required technical advancements progress in parallel.
@@
While this study provides valuable insights into SBSP integration within Europe’s energy system, several limitations should be acknowledged. First, the modeling of SBSP simplifies orbital energy dynamics by assuming a fixed shadow period and disregarding transient power fluctuations during Earth’s eclipse phases. Additionally, the study restricts SBSP deployment to equatorial GEO at 0° longitude, omitting considerations of alternative orbital placements or variations in altitude that could impact energy yield and transmission efficiency. Second, the construction of the 2050 energy system relies on 2020 climate data due to the absence of future climate data, which may not fully capture long-term meteorological trends affecting renewable generation. Third, the analysis does not assess SBSP’s role in grid stability, particularly its ability to provide ancillary services, manage fluctuations, or complement terrestrial renewables in real-time operations—an essential factor for large-scale integration. Fourth, the study does not account for geopolitical, regulatory, or spatial constraints that may impact SBSP deployment across different European nations, potentially limiting its feasibility in real-world implementation. Fifth, our SBSP performance and costs derive from NASA’s reports, and may not encompass the full spectrum of perspectives in the broader SBSP community, although our sensitivity analysis aims to mitigate this limitation.
@@
Looking ahead, future system-level research should expand to other regions beyond Europe, incorporate more precise satellite positioning and ground receiver optimization, address real-time operational feasibility (including ancillary services), and evaluate potential risks from heavy reliance on orbital infrastructure, such as space interference or equipment failure. Further exploration of geopolitical and policy factors—including collaborative frameworks for SBSP deployment—will also be essential to ensure resilient, equitable, and sustainable energy transitions.

### v2

@@ 3 Results
To assess the impact of SBSP integration, we select the Medium SBSP Market Opportunity Scenario as a representative case. Figure 4 compares the weekly average power outputs of SBSP, solar, wind, hydro, and storage technologies under this scenario. Throughout the year, heliostat design consistently delivers 300–350 GW with only minor dips around the spring and autumn equinoxes. By contrast, solar, wind, and hydro in both scenarios hover below 200 GW on average and exhibit notable seasonal swings—solar peaks in summer, while wind and hydro tend to peak in winter or other off-peak solar seasons.
@@
With heliostat design’s inclusion, wind generation declines by approximately 50%, while solar output drops by 70–80%, illustrating heliostat design’s displacement of intermittent renewables. In contrast, the planar design (RD2) is not selected as shown in Figure 4b (a no-SBSP case) in which the energy system relies more heavily on variable renewables throughout the year. This increases the need for storage to ensure supply-demand balance. The total storage usage displays larger capacity and wider fluctuations throughout the year than that in the scenario with RD1, underscoring the need for intensive storage interventions in the absence of SBSP.
@@
Figure 4: Weekly Power Output Variability of SBSP, Other Renewables and Storage Technologies (Discharging) under Medium SBSP Market Opportunity Scenario in 2050. Panels (a) and (b) present a comparative analysis of heliostat design (RD1) and planar design (RD2), respectively, with Solar, Wind, Hydro, Pumped Hydro Storage (PHS), Battery, and Hydrogen storage, in terms of power output and variability. Solid lines indicate weekly average power, while shaded bands denote the range of weekly maximum and minimum power fluctuations.
@@
The deployment of heliostat design leads to a substantial shift in storage utilization. In summer months, average power withdrawn from storage decreases by over 50% relative to the case without SBSP, primarily due to the consistent and high output from heliostat design reducing the need for balancing services. This is reflected in narrower deviation bands for all storage technologies, indicating a more consistent operating regime. In winter, however, total weekly storage requirements increase slightly, revealing a continued need for long-duration seasonal solutions. Although battery usage declines dramatically (by 78% annually), hydrogen storage takes on a larger role in winter, highlighting how heliostat design’s near-baseload profile mitigates short-term imbalances without fully eliminating seasonal challenges.
[…]
Cost thresholds for SBSP in the future (sensitivity analysis)
[…]
@@
Fig. 5 shows that once SBSP capital costs fall to around 14× or 9× the 2050 solar PV cost in heliostat design and planar design cases respectively, SBSP starts displacing wind and solar in a complementary role. A further approximately  35% cost reduction beyond these thresholds enables SBSP to emerge as the dominant baseload technology, surpassing conventional renewables in capacity deployment. However, at these levels, SBSP may produce surplus generation (due to reduced load factors, see SI Section 5), effectively meeting both baseload demand and balancing short- and long-duration fluctuations. However, current prohibitive costs of SBSP are at least 1-2 orders of magnitude higher than these complementary or baseload cost benchmarks.
@@
The results further highlight the shifting role of long-duration storage as SBSP scales from small to large deployment (Fig. 6). Initially, as SBSP capital costs decline to approximately 14× (heliostat design) or 9× (planar design) the projected 2050 solar PV cost, the capacity of short- and medium-duration storage technologies, such as batteries and PHS, steadily decreases. Hydrogen storage follows a similar trend, with a sharp reduction occurring when SBSP costs reach around 11–14× (heliostat design) or 8–9× (planar design) the 2050 solar PV benchmark. In this phase, SBSP effectively serves as a seasonal supplement, partially displacing the need for hydrogen storage.
@@
However, once SBSP capital costs fall below approximately 11× (heliostat design) or 8× (planar design) the 2050 solar PV cost, hydrogen storage capacity surges dramatically, reaching up to 178% (heliostat design) and 174% (planar design) of its original level in the absence of SBSP. This sudden expansion indicates that at these cost levels, winter electricity demand imbalances drive extensive reliance on hydrogen storage, despite SBSP’s ability to manage short-term fluctuations. Indeed, while total storage capacity increases during this phase, the actual energy supply from storage continues to decline (Fig.7). In this context, hydrogen provides up to 96% of its total annual supply during winter, highlighting the persistent need for long-duration storage to complement SBSP, even when SBSP reaches cost-competitive levels.
@@
A sharp decline in total storage capacity occurs when SBSP capital costs drop further to 8× (heliostat design) or 5× (planar design) the 2050 solar PV cost, marking the point at which SBSP surplus capacity becomes dominant. This shift signals SBSP’s ability to provide continuous baseload power, significantly reducing the system’s reliance on both short- and long-duration storage solutions.
@@
Figure 6: Impact of Heliostat Design (RD1) and Planar Design (RD2) Capital Cost on Optimized Energy Storage Capacity in 2050. The panel (a) and (b) illustrates the relationship between SBSP’s capital cost and the optimized energy storage capacity of three storage technologies: battery, hydrogen, and pumped hydro storage (PHS), as well as their total capacity, in the 2050 scenario. The energy storage capacity represents the maximum amount of electricity that storage devices can deliver when fully charged.
@@
Figure 7a and 7c illustrate the transition of SBSP into the dominant electricity source as its capital cost decreases. When SBSP capital costs reach approximately 9× (heliostat design) or 6× (planar design) the projected 2050 cost of solar PV, SBSP surpasses other technologies in electricity generation. A further decline to around 3× the solar PV cost results in heliostat design and planar design contributing 99% and 98% of total electricity generation, respectively.
@@
Figure 7: Impact of Heliostat Design (RD1) and Planar Design (RD2) Capital Cost on Annual Electricity Supply and System Costs in 2050. Panel (a) and (c) show the total annual electricity supply from all generation and storage technologies under varying capital costs of heliostat design and planar design, respectively. Numbers above each bar represent the combined electricity supply from the three storage technologies, and the black dashed line indicates the annual electricity demand. Panel (b) and (d) present the corresponding total annual system costs for each scenario. Numbers above each bar indicate the total annual costs under different SBSP capital cost assumptions.
Notably, heliostat design and planar design exhibit opposing impacts on energy storage requirements. As heliostat design deployment increases, total storage supply declines, shrinking to just 9% of its initial level. In contrast, planar design—due to its greater output variability—necessitates additional storage to compensate for fluctuations. Despite these divergent trends in overall storage supply, the share of battery consistently increases across both scenarios. When SBSP capital costs fall to 3× the cost of solar PV, electricity supply from battery accounts for 100% of total stored energy in both heliostat design and planar design cases. This finding underscores a critical threshold: if SBSP costs decline sufficiently, a power system without long-term storage could become viable.
Fig. 7b and 7d show that the total annual system cost declines significantly as SBSP capital costs decrease, once heliostat design or planar design is deployed. When their capital costs fall to 3× the projected 2050 solar PV cost, total system costs are reduced by 52% for heliostat design and 33% for planar design, underscoring SBSP’s substantial potential to lower overall energy expenditures. Throughout this transition, SBSP’s total generation cost initially rises due to increased deployment but subsequently declines sharply as generation costs decrease, ultimately displacing biomass and wind to become the dominant contributor to generation costs. Additionally, the reduction in SBSP generation costs leads to a corresponding decline in transmission costs, with DC transmission experiencing the most pronounced decrease. This finding further supports the conclusion from Fig. S8, which highlights SBSP’s superior ability to reduce long-distance DC transmission requirements.
@@ 4 Discussions
This study explores the system-wide impacts of two representative SBSP designs—heliostat design (RD1, near-constant, lower TRL) and planar design (RD2, partially intermittent, higher TRL)—under diverse European energy system conditions. When we apply NASA’s 2050 cost and performance assumptions, heliostat design consistently emerges as a cost-competitive option, capable of reducing total system costs by 7–15% relative to a no-SBSP scenario. Its near-baseload profile not only displaces a substantial share of wind and solar but also cuts battery storage usage by as much as 78%. At the same time, long-duration solutions, especially hydrogen, still expand in winter months, confirming SBSP’s ability to resolve short-term imbalances while underscoring ongoing seasonal needs. In contrast, planar design never appears in the least-cost mix at NASA’s 2050 cost projections, implying that its improved but still partially intermittent orbital geometry and higher capital expenses render it uncompetitive against both heliostat design and terrestrial renewables. Notably, the stable, cross-network generation offered by heliostat design’s power supply markedly reduces long-distance DC transmission by 31–70% across regions, suggesting that a centrally coordinated, high-availability resource can diminish reliance on interregional electricity flows.
@@
Our sensitivity analysis pinpoints the capital-cost thresholds at which each design moves from non-competitive to supplementary, and ultimately to a dominant baseload resource. At present, SBSP’s costs are 1–2 orders of magnitude above these break-even points. As heliostat design and planar design costs approach 9–14× that of 2050 solar PV, SBSP begins to complement terrestrial renewables; dropping further to around 6–9× enables a baseload role with surplus capacity for balancing. Below these thresholds, SBSP could start to become the primary generation source, though the system might still reply on energy storage. Conversely, planar design—given its partially intermittent orbit—would also require cost reductions to at least match the lower complementary threshold, plus a reliance on short-duration storage to handle variability.
@@
Collectively, our findings show that near-constant SBSP (as exemplified by heliostat design) may deliver significant system benefits, but also depends on technological breakthroughs (e.g., advanced orbital assembly) to overcome its currently low TRL. In the shorter term, less-intermittent planar design, with higher TRL, may make it an easier candidate for early demonstration, but only if its capital costs are reduced sufficiently to remain economically viable alongside terrestrial solar, wind, and energy storage. A coordinated development strategy could thus focus on planar design demonstrations first to refine core SBSP technologies—such as wireless power transmission and modular in-orbit assembly—while concurrently accelerating R&D for SBSP designs with more continuous power generation. Such a plan balances the practical realization of a tangible full-scale SBSP prototype with the longer-term potential for near-baseload SBSP to disrupt the broader renewables landscape once its more complex capabilities achieve higher maturity.
@@
To achieve the cost-competitiveness of SBSP, achieving a gigawatt-scale SBSP platform by the 2030s requires several key breakthrough advances. Although space-qualified photovoltaics can achieve high efficiency, deploying vast arrays—potentially on the order of several or tens of km² in a single platform—remains unproven. In-orbit manufacturing and autonomous assembly, both at low to mid TRL, need to advance in parallel with large-scale orbital robotics and modular construction. Equally crucial is wireless power transmission. While small-scale demonstrations, such as Caltech’s recent low-power orbit-to-Earth test [10], validate core concepts, maintaining and scaled-up operating highly focused microwave beams over tens of thousands of kilometers needs to be proven. Consequently, our capacity-expansion modeling, though primarily focused on the system-level economic and dispatch implications of SBSP, must treat the 2030 horizon as an early-stage full-scale demonstration phase, with true commercialization more plausibly materializing before 2050 if all required technical advancements progress in parallel.
@@
While this study provides valuable insights into SBSP integration within Europe’s energy system, several limitations should be acknowledged. First, the modeling of SBSP simplifies orbital energy dynamics by assuming a fixed shadow period and disregarding transient power fluctuations during Earth’s eclipse phases. Additionally, the study restricts SBSP deployment to equatorial GEO at 0° longitude, omitting considerations of alternative orbital placements or variations in altitude that could impact energy yield and transmission efficiency. Second, the modeling does not account for potential impacts from space-specific challenges such as orbital congestion, transmission interruptions, or beaming variability, which could influence SBSP reliability and operational performance. Further work will take these factors into consideration for improving the generation characteristics in the model. Third, the construction of the 2050 energy system relies on 2020 climate data due to the absence of future climate data, which may not fully capture long-term meteorological trends affecting renewable generation. Fourth, the analysis does not assess SBSP’s role in grid stability, particularly its ability to provide ancillary services, manage fluctuations, or complement terrestrial renewables in real-time operations—an essential factor for large-scale integration. Fifth, the study does not account for geopolitical, regulatory, or spatial constraints that may impact SBSP deployment across different European nations, potentially limiting its feasibility in real-world implementation. Sixth, our SBSP performance and costs derive from NASA’s reports, and may not encompass the full spectrum of perspectives in the broader SBSP community, although our sensitivity analysis aims to mitigate this limitation. In particular, advanced space power measurement techniques, such as interval-oriented eigensystem realization algorithms [47], reliability-constrained sliding mode control approaches [48], and multi-objective optimization-inspired regularization methods [49], are emerging as critical tools to better characterize SBSP performance in orbit.
Looking ahead, future SBSP research should prioritize the coordinated advancement of enabling technologies such as multi-junction and lightweight photovoltaic cells, modular in-orbit assembly methods, and wireless power transmission techniques, while also fostering close collaboration with policymakers to ensure alignment with emerging regulatory frameworks. Future studies should also focus on developing strategies to mitigate the risks posed by orbital debris to SBSP systems [50]. In addition, enhancing the resilience of SBSP system architectures and strengthening encryption protocols for secure energy transmission will be essential to protect against potential hostile actions. Improvements in module performance and cost-effectiveness, together with proactive policy engagement, will be critical to move SBSP from conceptual feasibility toward scalable implementation.
@@
At the system level, future research should expand beyond Europe to explore global SBSP deployment opportunities, incorporating more precise satellite positioning, ground receiver optimization, and real-time operational feasibility, including the provision of ancillary services. Future work should also explore the role of SBSP in supporting a more integrated energy system by incorporating cross-sectoral electrification and sector coupling, enabling a better assessment of SBSP’s contribution to system-wide flexibility and decarbonization. Additional efforts are needed to optimize hybrid configurations integrating multiple SBSP designs, assess the feasibility of intercontinental SBSP networks for regional balancing, and investigate social acceptance, electromagnetic safety, and regulatory challenges. Evaluating risks associated with heavy reliance on orbital infrastructure, such as space interference, system degradation, and equipment failure, alongside developing collaborative international frameworks, will be essential to ensure resilient, equitable, and sustainable energy transitions.
Beyond SBSP, some other emerging technologies promise firm, low-carbon power or help mitigate renewables’ variability. Advanced nuclear—encompassing small modular reactors (SMRs) and Gen IV designs, with fusion more distant—aims to deliver continuous, carbon-free electricity. Demonstration SMRs have entered operation or advanced construction in recent years [51], and a Gen IV pebble-bed reactor began supplying power in 2021 [52]. By contrast, SBSP is unlikely to mature before the 2040s, so advanced nuclear may achieve commercial readiness sooner.
For flexibility, low-carbon fuels and storage technologies can balance intermittent renewables over diverse timescales. A recent European pilot successfully ran a 12 MW turbine on 100% hydrogen in 2023 [53], illustrating the feasibility of hydrogen-based peaking or seasonal supply. Meanwhile, other long-duration storage solutions (flow batteries, compressed air, pumped hydro) can buffer multi-day shortfalls in wind or solar [54].
Taken together, these technologies offer a complementary portfolio. SBSP and advanced nuclear provide baseload-like generation, reducing storage needs; excluding firm low-carbon resources can drive steep cost increases near 100% renewables [55]. Although SBSP (low TRL, high uncertainty) and advanced nuclear (moderate TRL, complex licensing) each face distinct obstacles, integrating multiple firm and flexible options may yield a more resilient, cost-effective pathway to net zero in Europe. More research on the trade-off and uncertainty between these technologies is needed for insights of cost-effective, reliable net zero transition.